# An Invitation to Neuroalgebraic Geometry

### v1

@@ 4 Beyond the Algebraic
In this section, we discuss scenarios where it is possible to apply ideas from algebraic geometry to non-algebraic machine learning models.
@@ 4.2 Tropical Geometry
For MLPs with ReLU activation function – and, more generally, with piece-wise linear ones – the associated neuromanifold can be studied directly via an alternative version of algebraic geometry, i.e. tropical geometry (Maclagan & Sturmfels, 2021). The latter is based on an algebra where the standard operations of addition and multiplication in $\mathbb{R}$ are replaced by maximum (or minimum) and addition, respectively. This gives rise to a geometric theory which is ‘degenerate’, in a certain (precise) sense. Polynomial functions in this theory are continuous convex piece-wise linear functions. The connection between ReLU networks and tropical geometry has been established by Arora et al. (2016) and, more explicitly, by Zhang et al. (2018), where it is argued that any piece-wise linear function can be represented by such a network with enough layers.
Tropical geometry is closely related to combinatorics and, in particular, polyhedral geometry. This offers a plethora of tools from discrete mathematics to address geometric questions. In particular, the properties on the left-hand side of Table 1 can be studied by investigating the structure of combinatorial objects associated to neural networks, such as polytopes and fans (Montúfar et al., 2022; Brandenburg et al., 2024b). For example, recent works have explored these techniques to describe the fibers of the parameterization (Section 3.3) of ReLU networks (Brandenburg et al., 2024a), which relates to profound questions on decompositions of tropical rational functions (Tran & Wang, 2022).
@@ Acknowledgements
We thank Antonio Lerario for insightful discussions and Tomas Pajdla for helpful comments improving this manuscript. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.

### v2

@@ 5 Beyond the Algebraic
Throughout this work, we have focused on algebraic models. In this section, we discuss the possibility of applying similar ideas beyond the purely algebraic domain.
In our view, borrowing ideas from different fields of mathematics contributes to a holistic and multidisciplinary theory, which is essential for a fundamental understanding of deep learning. In fact, some mathematical ideas discussed in this work extend beyond the boundaries of algebraic geometry. For example, invariants such as dimension and covering number are defined for arbitrary metric spaces, singularities can be studied beyond algebraic varieties, the fiber-dimension theorem holds in great generality, and Morse theory is often formulated within the context of (smooth) manifolds. As a consequence, these general tools are applicable to non-algebraic models. Yet, they often result in a stronger incarnation when specialized to the algebraic context. For instance, the study of singularities is more tractable for algebraic models and, in addition to general Morse theory, algebraic geometry provides new concepts such as data discriminants and Euclidean distance degrees, enabling a more explicit understanding of loss landscapes.
In addition, we underline that ideas from algebraic geometry can be applied to non-algebraic models. Below, we discuss scenarios where such extensions of the methods of neuroalgebraic geometry are possible.
@@ 5.2 Tropical Geometry
For MLPs with ReLU activation function – and, more generally, with piece-wise linear ones – the associated neuromanifold can be studied directly via an alternative version of algebraic geometry, i.e. tropical geometry (Maclagan & Sturmfels, 2021). The latter is based on an algebra where the standard operations of addition and multiplication in $\mathbb{R}$ are replaced by maximum (or minimum) and addition, respectively. This gives rise to a geometric theory which is ‘degenerate’, in a certain (precise) sense. Polynomial functions in this theory are continuous convex piece-wise linear functions. The connection between ReLU networks and tropical geometry has been established by Arora et al. (2018a) and, more explicitly, by Zhang et al. (2018), where it is argued that any piece-wise linear function can be represented by such a network with enough layers.
Tropical geometry is closely related to combinatorics and, in particular, polyhedral geometry. This offers a plethora of tools from discrete mathematics to address geometric questions. In particular, the properties on the left-hand side of Table 1 can be studied by investigating the structure of combinatorial objects associated to neural networks, such as polytopes and fans (Montúfar et al., 2022; Brandenburg et al., 2024b). For example, recent works have explored these techniques to describe the fibers of the parameterization (Section 4.3) of ReLU networks (Brandenburg et al., 2024a), which relates to profound questions on decompositions of tropical rational functions (Tran & Wang, 2024).
@@ 6 Conclusions and Future Directions
Throughout this work, we have discussed several open problems and general questions in the research program of neuroalgebraic geometry. We conclude by briefly outlining a few concrete directions for the immediate future. A natural next step is to extend the algebro-geometric analysis to a broader class of neural architectures. For instance, skip connections are algebraic components widely deployed in modern machine learning. Incorporating skip connections might result in a smoothing of neuromanifold singularities (Orhan & Pitkow, 2018), making their analysis particularly appealing from the perspective of neuroalgebraic geometry. Similarly, it would be interesting to understand the neuroalgebraic geometry of architectures such as Graph Neural Networks (GNNs) and State Space Models (SSMs), which are popular in several domains, and are closely related to attention-based and convolutional networks. These models incorporate structures in data such as symmetries and recursion, falling into the scope of geometric and topological deep learning (see Section 2.1). Even further, understanding how structures in data can be reflected by geometric features of the neuromanifold represents a more general and fundamental question.
From a broader perspective, a desirable outcome of neuroalgebraic geometry is to suggest the development of novel models, as mentioned in Section 1. We believe that the power of algebraic geometry and its tools lies not only in providing a descriptive framework, but also in offering prescriptive principles for the design of theoretically-grounded models. Potentially, this could pave the way to new generations of neural architectures whose behavior is controlled by algebro-geometric invariants.
@@ Acknowledgements
We thank Antonio Lerario for insightful discussions and Tomas Pajdla for helpful comments improving an earlier version of this manuscript. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.